# Bigger is not Always Better: Scaling Properties of Latent Diffusion Models

### v1

@@ Abstract
Keywords:
Scaling Properties Latent Diffusion Models
@@ 1 Introduction
Latent diffusion models (LDMs) [55], and diffusion models in general, trained on large-scale, high-quality data [30, 62] have emerged as a powerful and robust framework for generating impressive results in a variety of tasks, including image synthesis and editing [55, 48, 9, 53, 49], video creation [42, 43, 70, 63], audio production [31], and 3D synthesis [28, 32]. Despite their versatility, the major barrier against wide deployment in real-world applications [12, 8] comes from their low sampling efficiency. The essence of this challenge lies in the inherent reliance of LDMs on multi-step sampling [66, 16] to produce high-quality outputs, where the total cost of sampling is the product of sampling steps and the cost of each step. Specifically, the go-to approach involves using the 50-step DDIM sampling [64, 55], a process that, despite ensuring output quality, still requires a relatively long latency for completion on modern mobile devices with post-quantization. In contrast to single shot generative models (e.g., generative-adversarial networks (GANs) [13]) which bypass the need for iterative refinement [13, 23], the operational latency of LDMs calls for a pressing need for efficiency optimization to further facilitate their practical applications.
@@
Recent advancements in this field [27, 74, 47, 24, 25, 8] have primarily focused on developing faster network architectures with comparable model size to reduce the inference time per step, along with innovations in improving sampling algorithms that allow for using less sampling steps [64, 11, 22, 34, 33, 71]. Further progress has been made through diffusion-distillation techniques [36, 59, 65, 61, 14, 41], which simplifies the process by learning multi-step sampling results in a single forward pass, and then broadcasts this single-step prediction multiple times. These distillation techniques leverage the redundant learning capability in LDMs, enabling the distilled models to assimilate additional distillation knowledge. Despite these efforts being made to improve diffusion models, the sampling efficiency of smaller, less redundant models has not received adequate attention. A significant barrier to this area of research is the scarcity of available modern accelerator clusters [20], as training high-quality text-to-image (T2I) LDMs from scratch is both time-consuming and expensive—often requiring several weeks and hundreds of thousands of dollars.
@@
In this paper, we empirically investigate the scaling properties of LDMs, with a particular focus on understanding how their scaling properties impact the sampling efficiency across various model sizes. We trained a suite of 12 text-to-image LDMs from scratch, ranging from 39 million to 5 billion parameters, under a constrained budget. Example results are depicted in Fig. 1. All models were trained on TPUv5 using internal data sources with about 600 million aesthetically-filtered text-to-image pairs. Our study reveals that there exist a scaling trend within LDMs, notably that smaller models may have the capability to surpass larger models under an equivalent sampling budget. Furthermore, we investigate how the size of pre-trained text-to-image LDMs affects their sampling efficiency across diverse downstream tasks, such as real-world super-resolution [58, 57] and subject-driven text-to-image synthesis (i.e., Dreambooth) [56].
@@ 1.1 Summary
Sampler does not change the scaling efficiency. Smaller models consistently demonstrate superior sampling efficiency, regardless of the diffusion sampler used. This holds true for deterministic DDIM [64], stochastic DDPM [16], and higher-order DPM-Solver++ [35]. For more details, see Section 3.4.
@@ 2 Related Work
Recent Large Language Models (LLMs) including GPT [3], PaLM [2], and LLaMa [67] have dominated language generative modeling tasks. The foundational works [21, 3, 18] for investigating their scaling behavior have shown the capability of predicting the performance from the model size. They also investigated the factors that affect the scaling properties of language models, including training compute, dataset size and quality, learning rate schedule, etc. Those experimental clues have effectively guided the later language model development, which have led to the emergence of several parameter-efficient LLMs [18, 67, 75, 1]. However, scaling generative text-to-image models are relatively unexplored, and existing efforts have only investigated the scaling properties on small datasets or small models, like scaling UNet [46] to 270 million parameters and DiT [47] on ImageNet (14 million), or less-efficient autoregressive models [6]. Different from these attempts, our work investigates the scaling properties by scaling down the efficient and capable diffusion models, i.e. LDMs [55], on internal data sources that have about 600 million aesthetics-filtered text-to-image pairs for featuring the sampling efficiency of scaled LDMs. We also scale LDMs on various scenarios such as finetuning LDMs on downstream tasks [69, 56] and distilling LDMs [41] for faster sampling to demonstrate the generalizability of the scaled sampling-efficiency.
@@
Nichol et al. [46] show that the generative performance of diffusion models improves as the model size increases. Based on this preliminary observation, the model size of widely used LDMs, e.g., Stable Diffusion [55], has been empirically increased to billions of parameters [51, 48]. However, such a large model makes it impossible to fit into the common inference budget of practical scenarios. Recent work on improving the sampling efficiency focus on improving network architectures [27, 74, 47, 24, 25, 8] or the sampling procedures [64, 11, 22, 34, 33, 71]. We explore sampling efficiency by training smaller, more compact LDMs. Our analysis involves scaling down the model size, training from scratch, and comparing performance at equivalent inference cost.
@@
Figure 1: Text-to-image results from our scaled LDMs (39M - 2B), highlighting the improvement in visual quality with increased model size (note: 39M model is the exception). All images generated using 50-step DDIM sampling. We use representative prompts from PartiPrompts [72], including “a professional photo of a sunset behind the grand canyon.”, “Dogs sitting around a poker table with beer bottles and chips. Their hands are holding cards.”, ‘Portrait of anime girl in mechanic armor in night Tokyo.”, “a teddy bear on a skateboard.”, “a pixel art corgi pizza.”, “Snow mountain and tree reflection in the lake.”, “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”, “a store front that has the word ‘LDMs’ written on it.”, and “ten red apples.”. Check our supplement for additional visual comparisons.
@@
Table 1: We scale the baseline LDM (i.e., 866M Stable Diffusion v1.5) by changing the base number of channels $c$ that controls the rest of the U-Net architecture as $[c,2c,4c,4c]$ (See Fig. 2). GFLOPS are measured for an input latent of shape $64\times 64\times 4$ with FP32. We also show a normalized running cost with respect to the baseline model. The text-to-image performance (FID and CLIP scores) for our scaled LDMs is evaluated on the COCO-2014 validation set with 30k samples, using 50-step DDIM sampling. It is worth noting that all the model sizes, and the training and the inference costs reported in this work only refer to the denoising UNet in the latent space, and do not include the 1.4B text encoder and the 250M latent encoder and decoder.
@@
Compared to diffusion models, other generative models such as, Variational Autoencoders (VAEs) [26, 54, 39, 68], Generative Adversarial Networks (GANs) [13, 40, 23, 52, 45], and Masked Models [10, 50, 15, 5, 4], are more efficient, as they rely less on an iterative refinement process. Sauer et al. [60] recently scaled up StyleGAN [23] into 1 billion parameters and demonstrated the single-step GANs’ effectiveness in modeling text-to-image generation. Chang et al. [4] scaled up masked transformer models for text-to-image generation. These non-diffusion generative models can generate high-quality images with less inference cost, which require fewer sampling steps than diffusion models and autoregressive models, but they need more parameters, i.e., 4 billion parameters.
@@ 3 Scaling LDMs
We developed a family of powerful Latent Diffusion Models (LDMs) built upon the widely-used 866M Stable Diffusion v1.5 standard [55] [^1] We adopted SD v1.5 since it is among the most popular diffusion models https://huggingface.co/models?sort=likes.. The denoising UNet of our models offers a flexible range of sizes, with parameters spanning from 39M to 5B. We incrementally increase the number of filters in the residual blocks while maintaining other architecture elements the same, enabling a predictably controlled scaling. Table 1 shows the architectural differences among our scaled models. We also provide the relative cost of each model against the baseline model. Fig. 2 shows the architectural differences during scaling. Models were trained using internal data sources with about 600 million aesthetically-filtered text-to-image pairs. All the models are trained for 500K steps, batch size 2048, and learning rate 1e-4. This allows for all the models to have reached a point where we observe diminishing returns. Fig. 1 demonstrates the consistent generation capabilities across our scaled models. We used the common practice of 50 sampling steps with the DDIM sampler, 7.5 classifier-free guidance rate, for text-to-image generation. The visual quality of the results exhibits a clear improvement as model size increases.
In order to evaluate the performance of the scaled models, we test the text-to-image performance of scaled models on the validation set of COCO 2014 [30] with 30k samples. For downstream performance, specifically real-world super-resolution, we test the performance of scaled models on the validation of DIV2K with 3k randomly cropped patches, which are degraded with the RealESRGAN degradation [69].
Refer to captionRefer to caption
Figure 3: In text-to-image generation, we observe consistent trends across various model sizes in how quality metrics (FID and CLIP scores) relate to training compute (i.e., the product of training steps and model GFLOPS). Under moderate training resources ($<1G$) training compute is the most relevant factor dominating quality. When training compute is unlimited (e.g., $\geq 1G$), larger models ($>700M$) lead to superior performance. Note that the 5B model’s size exceeds single TPU capacity, forcing memory-efficient training. This regime has different learning dynamics than smaller models (e.g., 2B), impacting accuracy at comparable training compute.
@@ 3.1 Training compute scales text-to-image performance
Refer to captionRefer to caption
Figure 4: In 4×4$\times$4 × real image super-resolution, FID and LPIPS scores reveal an interesting divergence. Model size drives FID improvement, while training compute most impacts LPIPS. Despite this, visual assessment (Fig. 5) confirms the importance of model size for superior detail recovery (similarly as observed in the text-to-image pretraining).
@@
Figure 5: In 4$\times$ super-resolution, visual quality directly improves with increased model size. As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities in both quantitative (Fig 4) and qualitative ways. Additional results are given in supplementary material.
@@
Figure 6: Visualization of the Dreambooth results shows two distinct tiers based on model size. Smaller models (83M-223M) perform similarly, as do larger ones (318M-2B), with a clear quality advantage for the larger group. Additional results are given in supplementary material.
@@ 3.2 Pretraining scales downstream performance
Using scaled models based on their pretraining on text-to-image data, we finetune these models on the downstream tasks of real-world super-resolution [58, 57] and DreamBooth [56]. The performance of these pretrained models is shown in Table. 1. In the left panel of Fig. 4, we present the generative performance FID versus training compute on the super-resolution (SR) task. It can be seen that the performance of SR models is more dependent on the model size than training compute. Our results demonstrate a clear limitation of smaller models: they cannot reach the same performance levels as larger models, regardless of training compute.
@@ 3.3 Scaling sampling-efficiency
Figure 7: Visualization of text-to-image results with different CFG rates (from left to right in each row: $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$). The prompt used is “A raccoon wearing formal clothes, wearing a top hat and holding a cane. Oil painting in the style of Rembrandt.”. We observe that changes in CFG rates impact visual quality more significantly than the prompt semantic accuracy. We use the FID score for quantitative determination of optimal sampling performance (Fig. 8) because it directly measures visual quality, unlike the CLIP score, which focuses on semantic similarity.
@@
Text-to-image generative models require nuanced evaluation beyond single metrics. Sampling parameters are vital for customization, with the Classifier-Free Guidance (CFG) rate [17] directly influencing the balance between visual fidelity and semantic alignment with text prompt. Rombach et al. [55] experimentally demonstrate that different CFG rates result in different CLIP and FID scores.
Refer to captionRefer to caption
Figure 9: Comparison of text-to-image performance of models with varying sizes. The left figure shows the relationship between sampling cost (normalized cost $\times$ sampling steps) and sampling steps for different model sizes. The right figure plots the text-to-image FID score as a function of the sampling cost for the same models. Key Observation: Smaller models achieve better FID scores than larger models for a fixed sampling cost. For instance, at a cost of 3, the 83M model achieves the best FID compared to the larger models. This suggests that smaller models might be more efficient in achieving good visual quality with lower costs. We highlight the best performing model at different sampling costs [3–48].

### v2

@@ 1 Introduction
Latent diffusion models (LDMs) (Rombach et al., 2022), and diffusion models in general, trained on large-scale, high-quality data (Lin et al., 2014; Schuhmann et al., 2022) have emerged as a powerful and robust framework for generating impressive results in a variety of tasks, including image synthesis and editing (Rombach et al., 2022; Podell et al., 2023; Delbracio & Milanfar, 2023; Ren et al., 2023; Qi et al., 2023), video creation (Mei & Patel, 2023; Mei et al., 2023; Wu et al., 2023; Singer et al., 2022), audio production (Liu et al., 2023a), and 3D synthesis (Lin et al., 2023; Liu et al., 2023b). Despite their versatility, the major barrier against wide deployment in real-world applications (Du et al., 2023; Choi et al., 2023) comes from their low sampling efficiency. The essence of this challenge lies in the inherent reliance of LDMs on multi-step sampling (Song et al., 2021b; Ho et al., 2020) to produce high-quality outputs, where the total cost of sampling is the product of sampling steps and the cost of each step. Specifically, the go-to approach involves using the 50-step DDIM sampling (Song et al., 2021a; Rombach et al., 2022), a process that, despite ensuring output quality, still requires a relatively long latency for completion on modern mobile devices with post-quantization. In contrast to single shot generative models (e.g., generative-adversarial networks (GANs) (Goodfellow et al., 2020)) which bypass the need for iterative refinement (Goodfellow et al., 2020; Karras et al., 2019), the operational latency of LDMs calls for a pressing need for efficiency optimization to further facilitate their practical applications.
@@
Recent advancements in this field (Li et al., 2023; Zhao et al., 2023; Peebles & Xie, 2023; Kim et al., 2023b; a; Choi et al., 2023) have primarily focused on developing faster network architectures with comparable model size to reduce the inference time per step, along with innovations in improving sampling algorithms that allow for using less sampling steps (Song et al., 2021a; Dockhorn et al., 2022; Karras et al., 2022; Lu et al., 2022a; Liu et al., 2023c; Xu et al., 2023). Further progress has been made through diffusion-distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023b; Gu et al., 2023; Mei et al., 2024a), which simplifies the process by learning multi-step sampling results in a single forward pass, and then broadcasts this single-step prediction multiple times. These distillation techniques leverage the redundant learning capability in LDMs, enabling the distilled models to assimilate additional distillation knowledge. Despite these efforts being made to improve diffusion models, the sampling efficiency of smaller, less redundant models has not received adequate attention. A significant barrier to this area of research is the scarcity of available modern accelerator clusters (Jouppi et al., 2023), as training high-quality text-to-image (T2I) LDMs from scratch is both time-consuming and expensive—often requiring several weeks and hundreds of thousands of dollars.
@@
In this paper, we empirically investigate the scaling properties of LDMs, with a particular focus on understanding how their scaling properties impact the sampling efficiency across various model sizes. We trained a suite of 12 text-to-image LDMs from scratch, ranging from 39 million to 5 billion parameters, under a constrained budget. Example results are depicted in Fig. 1. All models were trained on TPUv5 using internal data sources with about 600 million aesthetically-filtered text-to-image pairs. Our study reveals that there exist a scaling trend within LDMs, notably that smaller models may have the capability to surpass larger models under an equivalent sampling budget. Furthermore, we investigate how the size of pre-trained text-to-image LDMs affects their sampling efficiency across diverse downstream tasks, such as real-world super-resolution (Saharia et al., 2022; Sahak et al., 2023) and subject-driven text-to-image synthesis (i.e., Dreambooth) (Ruiz et al., 2023).
@@ 1.1 Summary
Sampler does not change the scaling efficiency. Smaller models consistently demonstrate superior sampling efficiency, regardless of the diffusion sampler used. This holds true for deterministic DDIM (Song et al., 2021a), stochastic DDPM (Ho et al., 2020), and higher-order DPM-Solver++ (Lu et al., 2022b). For more details, see Section 3.4.
@@ 2 Related Work
Recent Large Language Models (LLMs) including GPT (Brown et al., 2020), PaLM (Anil et al., 2023), and LLaMa (Touvron et al., 2023) have dominated language generative modeling tasks. The foundational works (Kaplan et al., 2020; Brown et al., 2020; Hoffmann et al., 2022) for investigating their scaling behavior have shown the capability of predicting the performance from the model size. They also investigated the factors that affect the scaling properties of language models, including training compute, dataset size and quality, learning rate schedule, etc. Those experimental clues have effectively guided the later language model development, which have led to the emergence of several parameter-efficient LLMs (Hoffmann et al., 2022; Touvron et al., 2023; Zhou et al., 2023; Alabdulmohsin et al., 2024). However, scaling generative text-to-image models are relatively unexplored, and existing efforts have only investigated the scaling properties on small datasets or small models, like scaling UNet (Nichol & Dhariwal, 2021) to 270 million parameters and DiT (Peebles & Xie, 2023) on ImageNet (14 million), or less-efficient autoregressive models (Chen et al., 2020). Different from these attempts, our work investigates the scaling properties by scaling down the efficient and capable diffusion models, i.e. LDMs (Rombach et al., 2022), on internal data sources that have about 600 million aesthetics-filtered text-to-image pairs for featuring the sampling efficiency of scaled LDMs. We also scale LDMs on various scenarios such as finetuning LDMs on downstream tasks (Wang et al., 2021; Ruiz et al., 2023) and distilling LDMs (Mei et al., 2024a) for faster sampling to demonstrate the generalizability of the scaled sampling-efficiency.
@@
Nichol et al. (Nichol & Dhariwal, 2021) show that the generative performance of diffusion models improves as the model size increases. Based on this preliminary observation, the model size of widely used LDMs, e.g., Stable Diffusion (Rombach et al., 2022), has been empirically increased to billions of parameters (Ramesh et al., 2022; Podell et al., 2023). However, such a large model makes it impossible to fit into the common inference budget of practical scenarios. Recent work on improving the sampling efficiency focus on improving network architectures (Li et al., 2023; Zhao et al., 2023; Peebles & Xie, 2023; Kim et al., 2023b; a; Choi et al., 2023; Mei et al., 2024b) or the sampling procedures (Song et al., 2021a; Dockhorn et al., 2022; Karras et al., 2022; Lu et al., 2022a; Liu et al., 2023c; Xu et al., 2023; Mei et al., 2025). We explore sampling efficiency by training smaller, more compact LDMs. Our analysis involves scaling down the model size, training from scratch, and comparing performance at equivalent inference cost.
@@
Figure 1: Text-to-image results from our scaled LDMs (39M - 2B), highlighting the improvement in visual quality with increased model size (note: 39M model is the exception). All images generated using 50-step DDIM sampling and CFG rate of 7.5. We use representative prompts from PartiPrompts Yu et al. (2022), including “a professional photo of a sunset behind the grand canyon.”, “Dogs sitting around a poker table with beer bottles and chips. Their hands are holding cards.”, ‘Portrait of anime girl in mechanic armor in night Tokyo.”, “a teddy bear on a skateboard.”, “a pixel art corgi pizza.”, “Snow mountain and tree reflection in the lake.”, “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”, “a store front that has the word ‘LDMs’ written on it.”, and “ten red apples.”. Check our supplement for additional visual comparisons.
@@
Table 1: We scale the baseline LDM (i.e., 866M Stable Diffusion v1.5) by changing the base number of channels $c$ that controls the rest of the U-Net architecture as $[c,2c,4c,4c]$ (See Fig. 2). GFLOPS are measured for an input latent of shape $64\times 64\times 4$ with FP32. We also show a normalized running cost with respect to the baseline model. The text-to-image performance (FID and CLIP scores) for all scaled LDMs is evaluated on the COCO-2014 validation set with 30k samples, using 50-step DDIM sampling and Classifier-free Guidance (CFG) with a rate of 7.5. It is worth noting that all the model sizes, and the training and the inference costs reported in this work only refer to the denoising UNet in the latent space, and do not include the 1.4B text encoder and the 250M latent encoder and decoder.
@@
Compared to diffusion models, other generative models such as, Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende & Mohamed, 2015; Makhzani et al., 2015; Vahdat & Kautz, 2020), Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Mao et al., 2017; Karras et al., 2019; Reed et al., 2016; Miyato et al., 2018), and Masked Models (Devlin et al., 2019; Raffel et al., 2020; He et al., 2022; Chang et al., 2022; 2023), are more efficient, as they rely less on an iterative refinement process. Sauer et al. (Sauer et al., 2023a) recently scaled up StyleGAN (Karras et al., 2019) into 1 billion parameters and demonstrated the single-step GANs’ effectiveness in modeling text-to-image generation. Chang et al. (Chang et al., 2023) scaled up masked transformer models for text-to-image generation. These non-diffusion generative models can generate high-quality images with less inference cost, which require fewer sampling steps than diffusion models and autoregressive models, but they need more parameters, i.e., 4 billion parameters.
Figure 3: In text-to-image generation using 50-step DDIM sampling and CFG rate of 7.5, we observe consistent trends across various model sizes in how quality metrics (FID and CLIP scores) relate to training compute (i.e., the total GFLOPS spend on training). Under moderate training resources, training compute is the most relevant factor dominating quality.
@@ 3 Scaling LDMs
We developed a family of powerful Latent Diffusion Models (LDMs) built upon the widely-used 866M Stable Diffusion v1.5 standard (Rombach et al., 2022) [^1] We adopted SD v1.5 since it is among the most popular diffusion models https://huggingface.co/models?sort=likes.. The denoising UNet of our models offers a flexible range of sizes, with parameters spanning from 39M to 5B. We incrementally increase the number of filters in the residual blocks while maintaining other architecture elements the same, enabling a predictably controlled scaling. Table 1 shows the architectural differences among our scaled models. We also provide the relative cost of each model against the baseline model. Fig. 2 shows the architectural differences during scaling. Models were trained using the web-scale aesthetically filtered text-to-image dataset, i.e., WebLI (Chen et al., 2022). All the models are trained for 500K steps, batch size 2048, and learning rate 1e-4. This allows for all the models to have reached a point where we observe diminishing returns. Fig. 1 demonstrates the consistent generation capabilities across our scaled models. We used the common practice of 50 sampling steps with the DDIM sampler, 7.5 classifier-free guidance rate, for text-to-image generation. The visual quality of the results exhibits a clear improvement as model size increases.
In order to evaluate the performance of the scaled models, we test the text-to-image performance of scaled models on the validation set of COCO 2014 (Lin et al., 2014) with 30k samples. For downstream performance, specifically real-world super-resolution, we test the performance of scaled models on the validation of DIV2K with 3k randomly cropped patches, which are degraded with the RealESRGAN degradation (Wang et al., 2021).
@@ 3.1 Training compute scales text-to-image performance
Figure 4: In 4×4$\times$4 × real image super-resolution using 50-step DDIM sampling, FID and LPIPS scores reveal an interesting divergence. Model size drives FID score improvement, while training compute most impacts LPIPS score. Despite this, visual assessment (Fig. 5) confirms the importance of model size for superior detail recovery (similarly as observed in the text-to-image pretraining).
@@
Figure 5: In 4$\times$ super-resolution using 50-step DDIM sampling, visual quality directly improves with increased model size. As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities in both quantitative (Fig 4) and qualitative ways. Additional results are given in supplementary material.
@@
Figure 6: Visualization of the Dreambooth results (using 50-step DDIM sampling and CFG rate of 7.5) shows two distinct tiers based on model size. Smaller models (83M-223M) perform similarly, as do larger ones (318M-2B), with a clear quality advantage for the larger group. Additional results are given in supplementary material.
@@ 3.2 Pretraining scales downstream performance
Using scaled models based on their pretraining on text-to-image data, we finetune these models on the downstream tasks of real-world super-resolution (Saharia et al., 2022; Sahak et al., 2023) and DreamBooth (Ruiz et al., 2023). The performance of these pretrained models is shown in Table. 1. In the left panel of Fig. 4, we present the generative performance FID versus training compute on the super-resolution (SR) task. It can be seen that the performance of SR models is more dependent on the model size than training compute. Our results demonstrate a clear limitation of smaller models: they cannot reach the same performance levels as larger models, regardless of training compute.
@@ 3.3 Scaling sampling-efficiency
Figure 7: Visualization of text-to-image results with 50-step DDIM sampling and different CFG rates (from left to right in each row: $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$). The prompt used is “A raccoon wearing formal clothes, wearing a top hat and holding a cane. Oil painting in the style of Rembrandt.”. We observe that changes in CFG rates impact visual quality more significantly than the prompt semantic accuracy. We use the FID score for quantitative determination of optimal sampling performance (Fig. 8) because it directly measures visual quality, unlike the CLIP score, which focuses on semantic similarity.
@@
Text-to-image generative models require nuanced evaluation beyond single metrics. Sampling parameters are vital for customization, with the Classifier-Free Guidance (CFG) rate (Ho & Salimans, 2022) directly influencing the balance between visual fidelity and semantic alignment with text prompt. Rombach et al. (Rombach et al., 2022) experimentally demonstrate that different CFG rates result in different CLIP and FID scores.
Figure 9: Comparison of text-to-image performance of models with varying sizes. The left figure shows the relationship between sampling cost (normalized cost $\times$ sampling steps) and sampling steps for different model sizes. The right figure plots the optimal text-to-image FID score among CFG rates of $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$ as a function of the sampling cost for the same models. Key Observation: Smaller models achieve better FID scores than larger models for a fixed sampling cost. For instance, at a cost of 3, the 83M model achieves the best FID compared to the larger models. This suggests that smaller models can be more efficient in achieving good results with lower costs.